DIRC Operations Report
Wednesday 2007/7/11 to Wednesday 2007/8/15
Oncall: Rolf Andreassen, Justine Serrano
Fast Monitoring: Rolf Andreassen, Justine Serrano, Jose Benitez
Luminosities this week
PEP peak lumi: ~1.08 x 1034
(last week: ~1.05 x 1034)
PEP delivery: 3.82 fb-1
(last weeks: 3.66 fb-1)
BaBar integrated lumi:3.68 fb-1
(last weeks: 3.21 fb-1)
Run 6 total delivery: 76.27 fb-1
Run 6 total BaBar recorded: 73.34 fb-1 total, including 6.21 fb-1 off-peak data
PEP activities/issues since last DIRC meeting
07/10: ROD. They fixed a vacuum leak that was causing a lot of background in the DIRC
07/17: A test to remove the abort gap in the beams was unsuccesful due to beam instabilities when filling the gap.
07/20: HER Beam-current monitor (BCM) broke so no trickling was possible until ~July 27. In the meantime the HER was operated on a fill-coast-fill mode.
08/01: Another leak in HER ring (IR2 7035) is causing vacuum spikes => Install helium lines to look for vacuum leak. Leak is far from the DIRC so background conditions are OK. Leak has not been found to this day.
08/02: No beam during most of the day to fix a water leak in the Linac which shorted a magnet.
08/03: Starting to have problem with the RF station 4-3, this is now limiting LER current. Possibly not fixable during this run, but they will replace it with a PS from the HER.
08/10: Switched to Offpeak running.
BaBar activities/issues since last DIRC meeting
07/17: DCH chiller broke. Karl replaced the filter and the pump. Didn't cause any down time.
08/11: BaBar magnet tripped causing a few hours downtime.
08/13: Since last week DCH has several LV drops in one FEA. They need a 1 shift long access to replace this FEA. This will probably happen next week.
DIRC activities/issues since last DIRC meeting
Integrated scaler saturation seen recently was due to a small gate in Jerry's electronics. Jerry fixed it.
07/19 : Matt disabled SOB dump valve for the UPS battery replacement, but the replacement did not happen, it will happen today.
07/31: Neutron counters' power supply tripped.
08/02: One christmas tree showed up in sector 5 during cosmic data taking. We turn off channel 6 in sector 5 (Set V0 and V1 at 110V) to avoid trip. There was another in sector 11, but it disappeared after turn off and on the high voltage (channels 5 and 10).
08/02: Masked out a noisy channel in sector 10.
08/03: We put scaler alarm thresholds back to their original values before July 11 vacuum leak.
08/06: Neutron counters' power supply tripped again, Matt replaced the power supply.
See DIRC logbook for details
Lumi and Background this month
- Take a standalone run at the next 1-2hour down time to investigate the Xmas trees.
- Replace four DFBs with goups of 8 dead or flaky channels:
- sector 0, DFB 0.
- 2 in sector 1, DFB 1 and DFB 13.
- sector 3, DFB 6.
- sector 8, DFB 0.
- Unplug christmas tree in sector 5 and sector 11.
- Replace air filters on Wienner crates.
- Investigate the crate of sector 8 which has again killed the first 8 channels of DFB 7.
| Equipment || # used || # spare (*) || Comments |
| DFB || 168 || 18
6 good (3F,4E,64,66,68,69),
3 potentially good (6E,57,9C)
9 bad (05,70,6C,78,40,88,95,87,5A) in France
| DCC || 12 || 7
3 good (12,31,5)
2 bad (25,9) in France
- Justine and Rolf rotating backup & on-call commissioning duties.
- I am leaving at the end of august, Loic will arrive in september.